# Enhancing Reasoning Skills in Small Persian Medical Language Models Can Outperform Large-Scale Data Training

### v3

@@ II-A Related Work In Medical Domain
One of the notable efforts in developing small-scale medical reasoning language models is MedSSS [10]. This framework focuses on fine-graining the intermediate reasoning steps to enhance the reasoning ability of medical models. To accomplish this, the authors applied the Monte Carlo Tree Search (MCTS) [11] algorithm on medical multiple-choice question-answering datasets to synthesize structured reasoning trajectories using a policy model. Based on this approach, they constructed three datasets supporting different stages of training: a dataset for Supervised Fine-Tuning (SFT) of the policy model, a rejected–preferred answer dataset for Direct Preference Optimization (DPO) ,and a soft dual-side label dataset for fine-tuning the Process Reward Model (PRM).Finally, the trained Policy Model was used as the primary reasoning policy, while the trained Process Reward Model acted as a decoding guide to refine and evaluate the reasoning process during generation.
Another significant contribution in this field is MedReason [12]. In this work, the authors leveraged a structured medical knowledge graph to transform conventional question–answer (Q&A) pairs into detailed medical reasoning trajectories. Each trajectory represented the step-by-step logical pathway connecting a clinical question to its correct answer, grounded in medical knowledge relationships such as symptoms, diagnoses, treatments, and physiological mechanisms. By constructing a reasoning-enriched dataset in this manner, the authors were able to fine-tune a baseline language model on these structured reasoning examples. This approach led to a marked improvement in the model’s ability to perform reasoning-intensive tasks within the medical domain, demonstrating the effectiveness of incorporating knowledge graph–based reasoning supervision into the training process. The findings from MedReason further emphasize the importance of structured reasoning representations as a means of improving the interpretability and analytical depth of medical language models.
Another important advancement in the development of medical reasoning models is HuatuoGPT-o1 [13]. In this work, the authors introduced a verification-guided reasoning framework designed to improve the logical consistency and accuracy of generated reasoning trajectories. Specifically, they employed a verifier model to assess and guide the policy model during reasoning generation, ensuring that each reasoning path adhered closely to factual correctness and medical plausibility. By filtering and refining reasoning trajectories through this verification process, they created a high-quality dataset composed of verified reasoning sequences. The authors then leveraged both supervised fine-tuning and reinforcement learning techniques to train their baseline language model using these verified data. This dual training approach allowed the model to not only imitate correct reasoning behaviors but also internalize reasoning principles through reward-driven optimization. As a result, HuatuoGPT-o1 demonstrated significant improvements in reasoning accuracy and reliability across a variety of medical question-answering and diagnostic tasks, highlighting the potential of verifier-guided learning frameworks in advancing medical reasoning language models.
@@ II-B Related Work In Other Domain
Perhaps the most influential recent work in the broader field of AI reasoning is DeepSeek-R [14]. Building upon the DeepSeek-V3-Base model, the authors introduced a reinforcement learning framework based on Group Relative Policy Optimization (GRPO) [15] to explicitly enhance the model’s reasoning capabilities. In this setup, the reinforcement learning process was guided by a reward function specifically designed to evaluate and maximize reasoning performance. Through this training paradigm, DeepSeek-R demonstrated remarkable improvements in logical reasoning and problem-solving accuracy across various benchmarks. However, the application of reinforcement learning also introduced several side effects. While reasoning performance improved substantially, the model’s readability and linguistic coherence deteriorated, and instances of language mixing became more frequent. To address these issues, the authors incorporated a small amount of cold-start supervised data and adopted a multi-stage training pipeline. This additional fine-tuning phase helped restore natural language fluency and readability while retaining the strong reasoning skills acquired through reinforcement learning. The resulting model, DeepSeek-R, thus represents a critical step forward in reasoning-oriented AI, demonstrating that reinforcement learning can significantly enhance reasoning ability—provided it is balanced with targeted fine-tuning to maintain linguistic quality.
Another notable contribution in the area of reasoning enhancement is Thought Preference Optimization (TPO) [16]. In this work, the authors proposed a preference-based framework for improving reasoning quality in language models. Given a question, the model first generates multiple candidate reasoning trajectories. These responses are then evaluated by a judge model, which identifies the best and worst samples based on reasoning correctness and coherence. The collected best–worst pairs are subsequently used to train the baseline model through Direct Preference Optimization (DPO), encouraging it to prefer higher-quality reasoning paths. This approach effectively aligns the model’s reasoning process with human-like evaluative feedback, demonstrating that reasoning quality can be substantially improved through structured preference optimization rather than relying solely on scale or supervised data.
Another relevant study was conducted by N. Ho et al. [17], who explored the transfer of reasoning capabilities from large language models to smaller ones. In their approach, a smaller model was fine-tuned using data generated by a larger model that exhibited stronger reasoning performance. The larger model produced reasoning trajectories and question–answer pairs that served as high-quality supervision signals for the smaller model. Through this distillation process, the smaller model effectively learned reasoning strategies and problem-solving patterns from its larger counterpart, achieving competitive reasoning performance with significantly reduced computational cost. This work demonstrates that reasoning ability can be efficiently transferred across models of different scales through targeted fine-tuning on reasoning-oriented synthetic data.
@@ III Proposed Methods
Two methods has been porposed. the first one, depicted in Figure 1, utilizes a teacher model (we used DeepSeek-R), which is itself a reasoning language model, to correct the student model’s medical reasoning mistakes in its response to the given multiple choice question. (if the student model chose the correct option we move to the next question in our dataset) When provided with the correct answer, the teacher generates a detailed chain-of-thought explanation that highlights the reasoning behind the solution. In this setup, the teacher’s output serves as the preferred answer, while the student’s original response is treated as the rejected answer for Direct Preference Optimization (DPO).
The second method, illustrated in Figure 2, employs the teacher model to provide more granular feedback by identifying specific errors in the student model’s answer to a given multiple-choice question. Upon receiving this feedback, the student model is prompted to correct its response precisely from the point where the error occurred. Through this iterative process, the student gradually refines its reasoning trajectory and converges on the correct answer. This constructive feedback allows the student model to develop reasoning independence, as it learns to correct its own mistakes rather than simply mimicking the teacher’s response.In this setup, the final answer generated by the student is considered the preferred response, while its initial attempt serves as the rejected one for DPO. Although this second framework is more time-consuming due to the iterative feedback process, it simplifies DPO training, since both the rejected and preferred answers are generated entirely by the student model itself as sequences of its own tokens.
Due to hardware limitations and the time efficiency of Framework 1, we trained the baseline model, aya-expanse-8b, using Framework 1 for 95% of the data and Framework 2 for the remaining 5%. This approach allowed us to efficiently train the model using Direct Preference Optimization (DPO). In total, the training involved 11,000 preferred–rejected answer pairs, containing approximately 2 million tokens in the preferred answers and 2.5 million tokens in the rejected ones.
@@ IV Data
IV-A PersianMedQA
Shortly after completing our translation process, a new Persian medical multiple-choice question answering dataset, called PersianMedQA, was introduced [22]. We randomly selected 1,000 questions from its training set and appended them to our dataset to incorporate additional data from a different source. The resulting dataset contained 19,000 Persian medical multiple-choice questions in total.
@@ VI Results
In the final evaluation, we compared the performance of gaokerena-V under direct (straight) prompting with that of gaokerena-R when provided with Chain-of-Thought (CoT) prompts. The results highlight that gaokerena-R, despite its smaller scale and limited training data, achieves superior reasoning performance through structured reasoning guidance, demonstrating the effectiveness of reasoning-centered optimization over pure data scaling.
@@ VI-C Final Evaluation
As previously discussed, gaokerena-V performs better when prompted directly, whereas gaokerena-R excels when prompted with a Chain-of-Thought (CoT) format that encourages step-by-step reasoning before producing an answer. One key advantage of CoT prompting—particularly when combined with multiple sampling and majority voting—is that it provides a natural measure of model certainty. If the generated responses converge on the same option across samples, the model can be considered confident; conversely, a wide dispersion among responses indicates uncertainty. This approach offers a practical alternative to directly querying the model’s self-assessed confidence [26] , a capability that smaller models generally lack due to their limited self-awareness and introspective reasoning abilities.
@@ VII Future Research
The results show that gaokerena-V performs better with direct prompts, while gaokerena-R excels with chain-of-thought (CoT) prompts. This indicates that both models are prompt-dependent, as their performance varies with the prompting style. Future research should aim to develop prompt-invariant medical language models that integrate strong reasoning skills and medical knowledge, achieving superior performance regardless of the prompt format. Achieving prompt invariance would represent an important step toward more reliable and generalizable small Persian medical language models.

### v4

@@ II-A Related Work In Medical Domain
The MedSSS framework is a notable, self-evolving system designed to instill robust, long-chain reasoning capabilities into small, deployable medical language models, such as the Llama3.1-8B-Instruct base model [10]. Its core methodology involves leveraging Monte Carlo Tree Search (MCTS) [11] over diverse medical datasets to construct rule-verifiable reasoning trajectories. These generated paths are used for policy refinement via supervised fine-tuning (SFT) and direct preference optimization (DPO), and to train a unique Process Reward Model (PRM). This PRM employs a soft dual-sided labeling objective that provides crucial step-level supervision, penalizing reasoning steps that degrade node value to mitigate hallucination and enhance interpretability in complex clinical reasoning. The resulting MedSSS system demonstrated state-of-the-art performance across eleven clinical reasoning benchmarks, achieving an average performance gain of +14.12 compared to its base model and significantly surpassing 32B-scale general-purpose reasoning models.
The MedReason [12] framework is another significant contribution that addresses the scarcity of high-quality medical reasoning data by leveraging a structured medical knowledge graph (KG). This approach converts conventional question–answer (Q&A) pairs into a dataset of 32,682 factually grounded reasoning trajectories, serving as structural supervision. Fine-tuning models on these paths consistently boosts performance: the framework achieved average accuracy gains of +5.4% for LLaMA 3.1-Instruct-8B and up to 7.7% for DeepSeek-Distill-8B. The resulting MedReason-8B model established a new state-of-the-art among 8B-parameter models, notably surpassing the Huatuo-o1-RL-8B by up to 4.2% on the MedBullets clinical benchmark, validating the importance of KG-based reasoning for interpretability and analytical depth.
The HuatuoGPT-o1 [13] framework is a two-stage, verification-guided system designed to instill complex reasoning in medical LLMs using 40K verifiable medical problems. Initially, a verifier guides search strategies to generate verified reasoning trajectories for Supervised Fine-Tuning (SFT), teaching the model to refine its answers. Subsequently, Reinforcement Learning (RL) is applied using verifier-based rewards for further enhancement. This methodology resulted in an average 8.5-point performance gain for the 8B model on medical benchmarks, with the complex Chain-of-Thought (CoT) strategy alone providing an average 4.3-point boost.
@@ II-B Related Work In Other Domain
The DeepSeek-R1 model [14] demonstrated that applying a reinforcement learning framework, specifically the resource-efficient Group Relative Policy Optimization (GRPO) [15] , to the DeepSeek-V3-Base model can effectively enhance complex reasoning capabilities. While the pure RL training (DeepSeek-R1-Zero) yielded powerful reasoning behaviors, it simultaneously introduced significant side effects, most notably a deterioration in readability and an increase in language mixing. To counteract these issues, the authors implemented a multi-stage curriculum that incorporated a cold-start supervised fine-tuning (SFT) phase. This targeted intervention proved essential for restoring the model’s linguistic quality and coherence. The resultant model, DeepSeek-R1, leveraged this carefully balanced training pipeline to achieve reasoning performance comparable to state-of-the-art models such as OpenAI-o1-1217, thereby confirming that the successful enhancement of reasoning through RL requires a subsequent, targeted fine-tuning stage to maintain linguistic integrity and coherence.
A notable contribution in the area of reasoning enhancement is Thought Preference Optimization (TPO), proposed by Wu et al [16]. as an iterative Reinforcement Learning from AI Feedback (RLAIF) approach. This preference-based framework compels the model to generate multiple candidate thought and response pairs, subsequently utilizing a judge model to evaluate only the resulting response quality. The collected best–worst pairs are then used to train the baseline model via Direct Preference Optimization (DPO), teaching it to generate high-quality internal thoughts without direct thought supervision. Utilizing an 8B parameter model as the base, this method demonstrated substantial performance gains in general instruction following, achieving an impressive 52.5% win rate on AlpacaEval (LC) and 37.3% on Arena-Hard. These results represent gains of over 4% compared to the direct response baseline, confirming that structured preference optimization effectively enhances reasoning capabilities, even across non-traditional domains such as marketing and health.
The study by N. Ho et al. [17] introduced Fine-tune-CoT, an effective knowledge distillation method that leverages prohibitively large language models (LLMs) to generate Chain-of-Thought (CoT) rationales for the fine-tuning of significantly smaller student models. This systematic transfer process enabled student models, which were approximately 25–100x smaller in parameter count, to acquire substantial reasoning capabilities. For instance, the 6.7B student model achieved an accuracy of 53.33% on MultiArith using diverse reasoning, demonstrating that complex, high-performance reasoning abilities can be efficiently transferred across a significant reduction in model scale.
@@ III Proposed Methods
We propose a two-stage framework to enhance the reasoning capabilities of small Persian medical language models. First, we employ Reinforcement Learning with AI Feedback (RLAIF) to construct a preference dataset $\mathcal{D}=\{(x,y_{w},y_{l})\}$, where $x$ is a medical question, $y_{w}$ is a preferred reasoning trajectory, and $y_{l}$ is a rejected one. Second, we fine-tune the student model $\pi_{\theta}$ using Direct Preference Optimization (DPO).
To generate high-quality preference pairs, we utilize a teacher-student architecture. The student ($\pi_{S}$) is the baseline Persian model, and the teacher ($\pi_{T}$) is a reasoning-capable model (DeepSeek-R). We employ two strategies to populate $\mathcal{D}$ based on the student’s performance:
  1. 1.
    Teacher Correction for Reasoning Alignment: This strategy accounts for 95% of our training data. The student model $\pi_{S}$ generates an initial response $y_{init}$ for a given question $x$. If the response is incorrect, $y_{init}$ is designated as the rejected response ($y_{l}$). The teacher model $\pi_{T}$ is then prompted with the ground truth answer and instructed to generate a correct, detailed Chain-of-Thought (CoT) explanation. This teacher-generated trajectory serves as the preferred response ($y_{w}$). The resulting preference pair $(x,y_{w},y_{l})$ explicitly contrasts flawed student reasoning with expert teacher reasoning. This approach is depicted in Figure 1.
  2. 2.
    Teacher-Guided Self-Correction: For the remaining 5% of the data, we employ an iterative feedback loop to encourage intrinsic error correction. This limited scale was primarily due to hardware constraints. If the student’s initial response $y_{init}$ is incorrect, the teacher generates a textual critique $c$ that identifies the error without revealing the correct option. The student is then prompted with $c$ to self-correct and generate a new response, $y_{retry}$. If $y_{retry}$ is correct, it is designated as the preferred response ($y_{w}$), and the original incorrect response $y_{init}$ remains the rejected response ($y_{l}$). This method ensures that the final $y_{w}$ is generated by the student model’s own distribution. The full process is detailed in Figure 2.
We optimize the student policy $\pi_{\theta}$ to satisfy the generated preferences using Direct Preference Optimization (DPO). DPO is an alignment method that avoids the need for a separate, expensive reward model. Instead, it optimizes the policy directly by minimizing the negative log-likelihood of the preferred response relative to the reference model $\pi_{ref}$ (the frozen aya-expanse-8b). The objective function is defined as:
$\begin{split}\mathcal{L}_{DPO}(\pi_{\theta};\pi_{ref})=&-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\\ &\hskip-20.00003pt\Biggl[\log\sigma\Biggl(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{ref}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{ref}(y_{l}|x)}\Biggr)\Biggr]\end{split}$
Where $\sigma$ is the sigmoid function and $\beta$ is a hyperparameter that controls the extent of the policy’s deviation from the reference model. This single objective implicitly maximizes the probability of the valid reasoning trajectory $y_{w}$ while suppressing the probability of the flawed reasoning path $y_{l}$. The final dataset comprised approximately 11,000 pairs, yielding about 2 million preferred tokens and 2.5 million rejected tokens.
@@ VI Results
In the final evaluation, we compared the performance of gaokerena-V under direct (straight) prompting with that of gaokerena-R when provided with Chain-of-Thought (CoT) prompts. The results highlight that gaokerena-R, despite its smaller scale and limited training data, achieves superior reasoning performance through structured reasoning guidance, demonstrating the effectiveness of reasoning-centered optimization over pure data scaling.  [^5] You can see some examples that show the enhancements in the Gaokerena-R model at github.com/Mehrdadghassabi/Gaokerena-R/tree/main/assets/examples
@@ VI-C Final Evaluation
As previously discussed, gaokerena-V performs better when prompted directly, whereas gaokerena-R excels when prompted with a Chain-of-Thought (CoT) format that encourages step-by-step reasoning before producing an answer. This observation indicates a high degree of prompt dependence, a key characteristic of both models in their current state.
One key advantage of CoT prompting—particularly when combined with multiple sampling and majority voting—is that it provides a natural measure of model certainty. If the generated responses converge on the same option across samples, the model can be considered confident; conversely, a wide dispersion among responses indicates uncertainty. This approach offers a practical alternative to directly querying the model’s self-assessed confidence [25] , a capability that smaller models generally lack due to their limited self-awareness and introspective reasoning abilities.
@@ VII Future Research
The results show that gaokerena-V performs better with direct prompts, while gaokerena-R excels with chain-of-thought (CoT) prompts. This indicates that both models are prompt-dependent, and future work must address this critical limitation.
Specifically, future research should aim to develop prompt-invariant medical language models that integrate strong reasoning skills and medical knowledge, achieving superior performance regardless of the prompt format. This involves resolving the inherent trade-off between the superior factual recall of a data-scaled model (like gaokerena-V) and the enhanced logical consistency of a reasoning-focused model (like gaokerena-R). Achieving prompt invariance would represent an important step toward more reliable and generalizable small Persian medical language models.